OpenStudy (ericthestein):

How can I find the coordinates of a point if I know the distance that point is to another point, and I know the coordinates of the second point mentioned?

OpenStudy (helpmeee123):

can i see?

OpenStudy (ericthestein):

Sure. One coordinate is (5,-1), and the distance from that to the unknown coordinate is 10.

OpenStudy (ericthestein):

Also the equation of the line is y = (4/3)x-7.66 if that helps

OpenStudy (welshfella):

the distance is sqrt( [( 5 - x)^2 + (-1-y)^2] where (x,y) is the unknown point AND THIS = 10 SO WE CAN WRITE (5 - X)^2 + (-1 - y)^2 = 100

OpenStudy (welshfella):

yes it does help - we couldnt find the coordinates without that. we need to solve the system of equations

OpenStudy (ericthestein):

Yeah, I tried setting the distance formula with those values equal to the distance (10) but didn't know where to go from there

OpenStudy (welshfella):

y = (4/3) x - 7.66 (5 - x)^2 + (-1-y)^2 = 100

OpenStudy (welshfella):

thats a bit messy use substitution for y from first equation into the second.

OpenStudy (welshfella):

I'd use wolframalpha to do it.

OpenStudy (ericthestein):

I may be wrong but so far I got 0=x^2-11.33x+68.34 The roots would be the possible x values right?

OpenStudy (welshfella):

there are 2 possible points

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

No there are no real roots for that one

OpenStudy (ericthestein):

Yeah. I did something wrong

OpenStudy (ericthestein):

Oh I just saw your wolframalpha link. Thanks

OpenStudy (ericthestein):

I see what I did wrong

OpenStudy (welshfella):

substituting ( 5 - x)^2 + (-1 - (4/3 x - 7.66)^2 = 100 (5 - x)^2 + (-(4/3)x + 6.66)^2 = 100 25 - 10x + x^2 + (16/9)x^2 - 24.54x + 44.3556) and so on..

OpenStudy (ericthestein):

Thanks for your time. I got it.

OpenStudy (welshfella):

yw